OpenStudy (anonymous):

Express the complex number in trigonometric form. -4i

OpenStudy (anonymous):

r(cos theta + i sin theta)

OpenStudy (ybarrap):

Use Euler's Identity: \(\large e^{i\theta} =\cos \theta+i\sin \theta\), where in your case \(\sin \theta =-4\) and \( \cos\theta=0\)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

i second that

OpenStudy (anonymous):



OpenStudy (anonymous):

r = sqrt(a^2 + b^2) theta = tan^(-1) (b/a)

OpenStudy (anonymous):

you need two numbers, \(r\) and \(\theta\) \(r\) is the absolute value, which is pretty clearly \(4\) as \(-4i\) is 4 units away from the origin

OpenStudy (anonymous):

and \(\theta\) you have a choice, but you can see the angle from the picture 

OpenStudy (anonymous):

you can use \(\theta=\frac{3\pi}{2}\) if you like or you could use \(\theta=-\frac{\pi}{2}\)